# Disambiguating Numeral Sequences to Decipher Ancient Accounting Corpora

### v1

@@ Abstract
1 Introduction
@@ 2 Data & Background
We base our analysis off the transliterated PE corpus hosted by the CDLI.111Cuneiform Digital Library Initiative, https://cdli.mpiwg-berlin.mpg.de; corpus downloaded 3 Oct 2022. Each text in this corpus contains an optional header, followed by a series of “entries” written one per line of the transliterated file. Each entry contains a (possibly empty) span of text, a comma delimiter, and a (possibly empty) numeral notation. The transliterations use a work-in-progress signlist that reflects experts’ current understanding of the texts, but which may not exactly match the true character inventory of the underlying script. In this signlist, characters that are believed to represent text are transliterated with labels beginning in “M” (e.g. M001, after Meriggi who pioneered the study of this script), and those representing digits are labeled with “N” (e.g. N01). The notation n(N00) means that the digit N00 is written down times. Figure 2 shows an example of a tablet alongside its transliteration.
@@
Most proto-Elamite numerals are written using one of four222Additionally, there are marginal systems (labeled B#, C#, and C”) which appear to be derived from one of the four main systems by the addition of hatch marks or boxes drawn around the digits. These systems are rare, and the extra hatching or boxing makes them trivial to identify, so we ignore them for the remainder of this work. number systems, which are called decimal (D), sexagesimal (S), bisexagesimal (B), and capacity (C). In spite of their names, all of these systems use mixed radices. Figure 1 shows the relative values of the digits in each of these systems, as derived through prior manual analyses Friberg (1978); Damerow and Englund (1989); Dahl (2019). Unlike Hindu-Arabic notation, where the value of a digit depends on its position, the values of proto-Elamite digits are fixed, and larger values are denoted by repeating a digit multiple times.
@@ 3.1 Automated Conversion
Of the 8011 intact numerals which we have extracted, there are 7954 for which this conversion returns at least one reading. Of these, only 1919 unambiguously belong to a particular number system: the remainder are ambiguous between two, three, or even all four systems (Table 1). The following sections outline two proposals for disambiguating the ambiguous cases. Section 4.1.1 discusses the 57 cases for which there is no valid reading in any system.
@@ 3.2 Subset-Sum Analysis
To achieve this, we filter the corpus to find texts with one or two entries on the reverse, as current understandings of the corpus suggest that these are likely to be summaries.333Some transliterations include an annotation which explicitly labels a particular entry as a summary. However, not all summaries are labeled in this way, so we rely on automatic detection of summaries to expand the number of texts available for this analysis. For each of these texts, we solve an instance of the subset-sum problem to identify whether any combination of readings from the obverse adds up to the same value as any reading of the reverse.
@@ 3.3 Bootstrapping
where is the set of features associated with sample , and is a learnable parameter which measures the association between feature and class . We apply the “cautious” approach from Collins and Singer 1999, which limits the number of rules that can be added to the decision list at each iteration of training. Specifically, candidate rules are sorted according to the number of labeled examples that support them, and only the with the largest support are added to the decision list.444Whitney and Sarkar 2012 note that many details are omitted from the description of the cautious algorithm in Collins and Singer 1999. We follow Whitney and Sarkar in assuming that confidence thresholding is performed using unsmoothed label counts. However, we differ from their approach by selecting the top rules overall (not the top for each label) as this yields stronger results on our data. starts at 5 and increases by 5 each iteration.
@@ 4.1.1 Invalid Notations
There are 57 intact numerals for which our automated conversion does not return a valid reading according to any number system. The vast majority of these violate the bundling principles established in prior work and shown in Figure 1. For example, the notation 11(N01) in P008043 should not be attested in any system, as every system carries over to a higher digit after at most ten N01s.
Some of these illegal notations may reflect errors on the part of the original scribes. For example, in P008844, the sum of the D values on the obverse equals 9(N23) 3(N14) 3(N01), or 573; the scribe has actually written 9(N23) 7(N14) 3(N01), which violates the usual principle that 6 N14s carry over to one N23. This suggests that the scribe may have conceived of the D system as a truly decimal notation (in which case the least significant digits would be written as 7 tens and 3 ones, exactly as we find on the tablet), forgetting that N23 uses a different radix. Several of the other aberrant notations lend credence to this view, such as P008788 which apparently records 88 M367s (likely goats, usually counted with D) as 8(N14) 8(N01). These cases suggest a lack of standardisation across scribes or across documents, which is consistent with the longstanding view that the writing system never achieved a significant degree of standardisation Dahl (2019).
@@ 4.3 Bootstrapping
Figure 3 and Table 5 compare results from our two approaches to bootstrapping. The baseline, vanilla bootstrapping algorithm achieves a mediocre 0.60 F1. Recall of the B and S classes is perfect, but only half of the instances of the C class are correctly labeled. This classifier does not seem to have uncovered any clear signals to identify the D system: 3 instances of this class remain unlabeled (none of their associated features ever attained the required confidence to be added to the decision list), and those which are labeled are distributed uniformly across all of the possible classes.
@@
| Baseline | 0.64 | 0.56 | 0.60 | 0.74 | 0.65 | 0.69 |
| Ours | 0.88 | 0.88 | 0.88 | 0.90 | 0.90 | 0.90 |
By contrast, our proposed approach to cautious rule selection yields 0.88 F1, with significantly better recall of both the C and D classes than the baseline. This suggests that frequency-based caution may be ill-suited for bootstrapping on some datasets, and that, in settings where bootstrap classifiers remain viable, it may be worthwhile to explore alternative approaches to cautious rule selection.
@@
Thus many numerals which are technically ambiguous between these systems nonetheless have unambiguous values.555Here we assume that the absolute value of N01 is the same across all three systems, and not just its value relative to the other digits. Current understandings suggest that this is the case, but the undeciphered nature of the script means this is technically not certain.
@@
In this two-way setting, our model achieves 0.90 F1, versus 0.69 F1 from the vanilla baseline (Figure 4).
We emphasize that our test set only contains numerals which domain experts were able to disambiguate based on manual inspection. Easy cases may therefore be over-represented, and we expect our results to be an upper bound on the classifier’s accuracy across the whole corpus. Despite this, the results are strong enough to suggest that a majority of the ambiguous numerals in the corpus can be disambiguated with some certainty.
@@ 5 Analysis
M376 has been suggested as either a "high-status human" (Dahl, 2005, p. 95) or livestock (Kelley, 2018, p. 165); it is strongly predicative of the sexigesimal system in our analysis. Other objects associated with this system include M056f (a plow), M219, M371 and M296 (very speculatively, syllables used to write personal names), M059, M145, and M365 (“owners”, possibly persons or institutions to whom these accounts belonged), and M269c (a vessel?). This may cast some doubt on the livestock reading for M376, in favor of the high-status human reading which is a more natural fit among the owner signs and possible personal names in this collection.
We note one text, P008212, which may exhibit a consistent ratio related to M376. P008212 alternates between entries ending in M288 and entries ending in M376 or M367i. The magnitude of the numeral in an M288 entry is always exactly 4 times as large as the magnitude of an adjacent M376 entry, or 2 times as large as an adjacent M367i entry. This pattern is very unlikely to be due to chance, as it holds across 44 total entries. Whatever the meaning of M376, on the basis of this text we can assert that it is associated with amounts of M288 that are exactly twice as large as those associated with M367i. To our knowledge this ratio has not yet been noted in previous publications.
@@
The numerical systems of proto-Elamite have been proposed to have functional uses relating to cultural practices in 3rd millennium south-western Iran. For instance, the capacity system (C) is suggested to be used for counting rations disbursed to households or workers (Kelley, 2018, pgs. 153-155). Among the recipients of these rations are M388 and M124, parallel “worker categories” which may represent the heads of work teams. We find that entries beginning in M388 accompany significantly larger capacity measures on average than those in M124, which may point towards M388 individuals heading teams of larger sizes or comprising workers of higher status.
@@ A.3 Decimal Measures
This tablet is broken: a fracture on the obverse renders one digit partially unreadable. The broken digit has been restored as 2(N23), but we propose that the correct reading should in fact be 1(N23).
To see that this should be the case, notice that the first entry on the reverse (P008179:14:num) unambiguously equals N01D; this is exactly the same sum obtained by reading the obverse entries in the decimal system, provided one uses our restoration instead of the current transliteration.
This reading is consistent with the fact that some entries (P008179:8:num, P008179:9:num, P008179:14:num) are already unambiguous decimal counts. Moreover, all but one of the entries on the obverse count the same object (M388), which further points towards their likely using the same number system. On the basis of this evidence, we disambiguate P008179:10 and P008179:11 to the decimal system.
The original restoration (2(N23)) presumably arose from the observation that the lacuna is wide enough to span two signs. The last visible sign before the lacuna is M388, which only occurs at the very end of entries elsewhere in this text, implying that all of the missing signs are likely digits. Our proposed reading requires that the N23 in the lacuna occupy a slightly wider space than would be typical, or perhaps that some of the lacuna be occupied by an erasure. If the original restoration is correct, there must instead be an arithmetic error in the summary line (which in this case differs from the true sum by 1(N23)).

### v2

@@ Abstract
1 Introduction111Englund 1996 contains a printing error in the table which reports the relative values of signs in the decimal system. An earlier version of this paper relied on the erroneous values from that publication. This is an updated version which corrects those mistakes.
@@ 2 Data & Background
We base our analysis off the transliterated PE corpus hosted by the CDLI.333Cuneiform Digital Library Initiative, https://cdli.mpiwg-berlin.mpg.de; corpus downloaded 3 Oct 2022. Each text in this corpus contains an optional header, followed by a series of “entries” written one per line of the transliterated file. Each entry contains a (possibly empty) span of text, a comma delimiter, and a (possibly empty) numeral notation. The transliterations use a work-in-progress signlist that reflects experts’ current understanding of the texts, but which may not exactly match the true character inventory of the underlying script. In this signlist, characters that are believed to represent text are transliterated with labels beginning in “M” (e.g. M001, after Meriggi who pioneered the study of this script), and those representing digits are labeled with “N” (e.g. N01). The notation n(N00) means that the digit N00 is written down times. Figure 2 shows an example of a tablet alongside its transliteration.
@@
Most proto-Elamite numerals are written using one of four444Additionally, there are marginal systems (labeled B#, C#, and C”) which appear to be derived from one of the four main systems by the addition of hatch marks or boxes drawn around the digits. These systems are rare, and the extra hatching or boxing makes them trivial to identify, so we ignore them for the remainder of this work. number systems, which are called decimal (D), sexagesimal (S), bisexagesimal (B), and capacity (C). In spite of their names, most of these systems use mixed radices. Figure 1 shows the relative values of the digits in each of these systems, as derived through prior manual analyses Friberg (1978); Damerow and Englund (1989); Dahl (2019). Unlike Hindu-Arabic notation, where the value of a digit depends on its position, the values of proto-Elamite digits are fixed, and larger values are denoted by repeating a digit multiple times.
@@ 3.1 Automated Conversion
Of the 8011 intact numerals which we have extracted, there are 7984 for which this conversion returns at least one reading. Of these, only 1899 unambiguously belong to a particular number system: the remainder are ambiguous between two, three, or even all four systems (Table 1). The following sections outline two proposals for disambiguating the ambiguous cases. Section 4.1.1 discusses the 27 cases for which there is no valid reading in any system.
@@ 3.2 Subset-Sum Analysis
To achieve this, we filter the corpus to find texts with one or two entries on the reverse, as current understandings of the corpus suggest that these are likely to be summaries.555Some transliterations include an annotation which explicitly labels a particular entry as a summary. However, not all summaries are labeled in this way, so we rely on automatic detection of summaries to expand the number of texts available for this analysis. For each of these texts, we solve an instance of the subset-sum problem to identify whether any combination of readings from the obverse adds up to the same value as any reading of the reverse.
@@ 3.3 Bootstrapping
where is the set of features associated with sample , and is a learnable parameter which measures the association between feature and class . We apply the “cautious” approach from Collins and Singer 1999, which limits the number of rules that can be added to the decision list at each iteration of training. Specifically, candidate rules are sorted according to the number of labeled examples that support them, and only the with the largest support are added to the decision list.666Whitney and Sarkar 2012 note that many details are omitted from the description of the cautious algorithm in Collins and Singer 1999. We follow Whitney and Sarkar in assuming that confidence thresholding is performed using unsmoothed label counts. However, we differ from their approach by selecting the top rules overall (not the top for each label) as this yields stronger results on our data. starts at 5 and increases by 5 each iteration.
@@ 4.1.1 Invalid Notations
There are 27 intact numerals for which our automated conversion does not return a valid reading according to any number system. The vast majority of these violate the bundling principles established in prior work and shown in Figure 1. For example, the notation 11(N01) in P008043 should not be attested in any system, as every system carries over to a higher digit after at most ten N01s. These cases may be simple scribal errors, or they may suggest a lack of standardisation across scribes or across documents, which is consistent with the longstanding view that the writing system never achieved a significant degree of standardisation Dahl (2019).
More rarely, these conversion failures result from the presence of sign names such as N02 which are not part of any PE number system established in prior work. Often these aberrant sign names correspond to digits from the related proto-cuneiform script. The number systems of proto-Elamite and proto-cuneiform bear a close resemblance, and in short tablets with minimal text content it can be difficult to confidently assign a text to one script or the other. For these texts, expert evaluation may be needed to establish whether the aberrant sign names reflect simple transliteration errors or the erroneous inclusion of proto-cuneiform material in the proto-Elamite corpus.
@@ 4.3 Bootstrapping
Figure 3 and Table 5 compare results from our two approaches to bootstrapping. The baseline, vanilla bootstrapping algorithm achieves a reasonable F1 of 0.88. Recall of the B, D, and S classes is strong, but the model performs more poorly on C notations, assigning these to the wrong system in of cases.
@@
| Baseline | 0.88 | 0.88 | 0.88 | 0.90 | 0.90 | 0.90 |
| Ours | 0.94 | 0.94 | 0.94 | 0.96 | 0.96 | 0.96 |
By contrast, our proposed approach to cautious rule selection yields 0.94 F1, and completely eliminates the confusion between the C and B systems. This suggests that frequency-based caution may not be the most appropriate rule-selection heuristic for all datasets, and that, in settings where bootstrap classifiers remain viable, it may be worthwhile to explore alternative approaches to cautious rule selection. We also note that our approach converges to a solution extremely quickly, completing after just 25 iterations of rule updates, versus more than 200 iterations for the baseline model.
@@
Thus many numerals which are technically ambiguous between these systems nonetheless have unambiguous values.777Here we assume that the absolute value of N01 is the same across all three systems, and not just its value relative to the other digits. Current understandings suggest that this is the case, but the undeciphered nature of the script means this is technically not certain.
@@
In this two-way setting, our model achieves 0.96 F1, versus 0.90 F1 from the vanilla baseline (Figure 4).
We emphasize that our test set only contains numerals which domain experts were able to disambiguate based on manual inspection. Easy cases may therefore be over-represented, and we expect our results to be an upper bound on these classifiers’ accuracy across the whole corpus. Despite this, the results are strong enough to suggest that a majority of the ambiguous numerals in the corpus can be disambiguated with some certainty.
@@ 5 Analysis
M376 has been suggested as either a "high-status human" (Dahl, 2005, p. 95) or livestock (Kelley, 2018, p. 165); it is strongly predicative of the sexigesimal system in our analysis. Other objects associated with this system include M056f (a plow), M219, M387c and M320 (very speculatively, syllables used to write personal names), M059, M145, M153, and M365 (“owners”, possibly persons or institutions to whom these accounts belonged), and M269c (a vessel?). This may cast some doubt on the livestock reading for M376, in favor of the high-status human reading which is a more natural fit among the owner signs and possible personal names in this collection.
We note one text, P009343,888An earlier version of this work mistakenly cited this text as P008212. which may exhibit a consistent ratio related to M376. P009343 alternates between entries ending in M288 and entries ending in M376 or M367i. The magnitude of the numeral in an M288 entry is always exactly 4 times as large as the magnitude of an adjacent M376 entry, or 2 times as large as an adjacent M367i entry. This pattern is very unlikely to be due to chance, as it holds across 44 total entries. Whatever the meaning of M376, on the basis of this text we can assert that it is associated with amounts of M288 that are exactly twice as large as those associated with M367i. To our knowledge this ratio has not yet been noted in previous publications.
@@
The numerical systems of proto-Elamite have been proposed to have functional uses relating to cultural practices in 3rd millennium south-western Iran. For instance, the capacity system (C) is suggested to be used for counting rations disbursed to households or workers (Kelley, 2018, pgs. 153-155). Among the recipients of these rations are M388 and M124, parallel “worker categories” which may represent the heads of work teams. We find that tablets whose first entry ends in M388 contain significantly larger capacity measures on average than those ending M124, which may point towards M388 individuals heading teams of larger sizes or comprising workers of higher status.
@@ A.3 Decimal Measures
This tablet contains an unambiguous summary entry ending in M388 with a value of N01D. This is exactly the value of the entries on the obverse (which also end in M388) when read in the decimal system.